# An upper bound of the silhouette validation metric for clustering

### v1

@@ 4.2 Derivation
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the other data points in the same cluster is not smaller than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning via the $\Lambda$-quotient of point $i$.
@@
for some $\kappa\in\{2,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
@@ 20  Append 1q1-q to AA;
The workflow of the proposed upper bound on the ASW is shown in Fig.8. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the (restricted) solution space $\min_{I}|C_{I}|\geq\kappa$. The steps are as follows: (1) For each row $i$ of $\Delta$, we sort the off-diagonal entries increasingly to obtain $\hat{\Delta}_{i,1},\hat{\Delta}_{i,2},\ldots$, the nearest-to-farthest distances from $i$. (2) For $i$, we initialize the running sums $x$ and $y$ so that $x/(\kappa-1)$ is its mean distance to its $\kappa-1$ nearest points, and $y/(n-\kappa)$ is its mean distance to its $n-\kappa$ farthest points. (3) Starting from $\Lambda=\kappa+1$, we update these running sums as $\Lambda$ increases, thereby scanning all admissible values $\Lambda\in\{\kappa,\ldots,n-\kappa\}$. For each $\Lambda$, we form the $\Lambda$-quotient $q(i,\Delta,\Lambda)$ and keep the minimum. (4) Finally, we record 1q1-q as the upper bound for $s(i\,|\,\mathcal{C}_{K},\Delta)$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound in Eq. (1) when $\kappa=1$.
Complexity
Forming $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$; $\mathcal{O}(n^{2})$ [54].
@@
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). To reflect the spread of the individual bounds, we denote by $\texttt{minUB}(\Delta)$ and $\texttt{maxUB}(\Delta)$ the minimum and maximum values of $A$, respectively. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space.
@@
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$; its tightness is data-dependent. Nonetheless, it provides an enlightening piece of information that merits inclusion in analyses.
4.4 Example
Recall the toy data from Tab. 3. We have $n=5$, so $\Lambda\in\{1,2,3,4\}$. The values $q(i,\Delta,\Lambda)$ for all instances are shown in Tab. 5. Consider the K=2K{=}2 clustering from Fig. 5: $C_{1}=\{x_{1},x_{2},x_{3}\}$ and $C_{2}=\{x_{4},x_{5}\}$. Tab. 6 compares each silhouette width with the corresponding upper bound, illustrating that $s(i)\leq 1-f(i,\Delta)$ for all $i$, with equality for $x_{2},x_{4},x_{5}$ and a small strict headroom for $x_{1}$ ($0.026$) and $x_{3}$ ($0.054$). Moreover, we have
$\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)=0.7512\;\leq\;0.7672=1-\frac{1}{n}\sum_{i=1}^{n}f(i,\Delta).$
Table 5: All possible values of $q(i,\Delta,\Lambda)$ for each point $i$ in the toy dataset.
Point $q(\cdot,1)$ $q(\cdot,2)$ $q(\cdot,3)$ $q(\cdot,4)$
$x_{1}$ 1.000 0.184 0.210 0.466
$x_{2}$ 1.000 0.417 0.348 0.531
$x_{3}$ 1.000 0.199 0.253 0.502
$x_{4}$ 1.000 0.220 0.538 0.643
$x_{5}$ 1.000 0.213 0.568 0.654
Table 6: Showing $s(i)\leq 1-f(i,\Delta)$ for each point in the toy dataset.
$i$ $|C_{I}|$ $\Lambda^{\star}=\arg\min_{\Lambda}q(i,\Delta,\Lambda)$ $f(i,\Delta)$ $1-f(i,\Delta)$ $s(i)$
$x_{1}$ 3 2 0.184 0.816 0.790
$x_{2}$ 3 3 0.348 0.652 0.652
$x_{3}$ 3 2 0.199 0.801 0.747
$x_{4}$ 2 2 0.220 0.780 0.780
$x_{5}$ 2 2 0.213 0.787 0.787
Figure 8: Workflow of the proposed upper bound on the $\mathrm{ASW}$.
4.5 Enabling early stopping of optimization loops through the upper bound
We now demonstrate how the data–dependent upper bound introduced in Eq. (1) can be used to accelerate standard ASW based optimization loops. In many practical applications, selecting the number of clusters involves iterating over $K$ and evaluating the output of the clustering algorithm $\mathcal{A}(K)$ in order to maximize the ASW. When the dataset is large, each iteration can be computationally demanding, and exhaustive searches over all candidate $K$ quickly become demanding. By incorporating the upper bound into this procedure, we can enable early stopping: if the current ASW is provably within a user–specified relative tolerance of the global optimum, the loop halts, saving potentially substantial computation while still guaranteeing near–optimal clustering performance within the search space. This type of procedure is described in more detail in Algo. 2.
Input: $\Delta\in\mathbb{R}^{n\times n}$;
  // Dissimilarity matrix
1
Input: $\epsilon$;
  // Relative error tolerance, $0\leq\epsilon\leq 1$
2
Input: $\tau$;
  // Clusterability threshold, $0\leq\tau\leq 1$
3
Input: $\mathcal{A}(K),K_{\max}$;
  // Clustering algorithm, max. no. clusters
4
Output: ASW-optimal clustering of $\Delta$ (within search space)
Compute $\mathrm{UB}(\Delta)$;
  // Apply Algo. 1
5 if $\mathrm{UB}(\Delta)\leq\tau$ then // Abort if dataset is insufficiently clusterable
6 return NULL
7optimal clustering $\leftarrow$ NULL ;
8 optimal ASW $\leftarrow$ -1 ;
9 for $K=2,\ldots,K_{\max}$ do
[…]
11      optimal clustering $\leftarrow\mathcal{A}(K)$ ;
12      optimal ASW $\leftarrow\mathrm{ASW}(\mathcal{A}(K),\Delta)$ ;
13    if $\frac{\mathrm{UB}(\Delta)-\text{optimal ASW}}{\mathrm{UB}(\Delta)}<\epsilon$ then
14       return optimal clustering
15
16
return optimal clustering
Algorithm 2 Early stopping in ASW-based optimization
Workflow
The procedure integrates the proposed upper bound into a standard ASW optimization loop. At each step, the current best clustering is compared to the global upper bound. If the relative gap falls below the user-specified tolerance $\epsilon$, the algorithm terminates early, certifying that no further search can improve ASW by more than $\epsilon\%$ of the true global optimum. In the worst case, the condition is never met, and the loop reverts to the standard exhaustive search over $K=2,\ldots,K_{\max}$, ensuring no loss in clustering quality.
Remark
If minimal cluster size constraints are of interest, the procedure can be modified by replacing the initial call to $\mathrm{UB}(\Delta)$ with the restricted version $\mathrm{UB}_{\kappa}(\Delta)$. This enables early stopping under user-specified requirements. We also emphasise that clusterability in this context is interpreted solely through the lens of the silhouette coefficient. While this provides a useful and practical criterion for our purposes, it is by no means a universal or definitive notion of clusterability—many alternative perspectives exist in the literature.
Recall the toy dataset summarized in Tab. 4. In Sec.  4.4, we established that the corresponding upper bound is $\texttt{UB}(\Delta)=0.7672$. Let $\widehat{S}_{K}$ denote the empirically observed $\mathrm{ASW}$ value for $K$ clusters. For the toy dataset, these values are:
$\widehat{S}_{2}=0.7512,\quad\widehat{S}_{3}=0.5173,\quad\widehat{S}_{4}=0.3068,\quad\widehat{S}_{5}=0.0000.$
The performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes the clustering achieving the global maximum ASW. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}_{K}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\widehat{S}_{K}$ is guaranteed to be within $\epsilon$ of the true maximum ASW.
Table 7: Early stopping on the toy dataset.
$K$ $\widehat{S}_{K}$ $\texttt{UB}(\Delta)$ worst-case relative error
[…]
As illustrated by Tab. 7, applying Algo. 2 with relative error tolerance $\epsilon=0.05$ for the toy dataset would halt the process already at $K=2$, thereby avoiding unnecessary evaluations for $K=3,4,5$. In this way, the upper bound allows substantial savings while still certifying near-optimal performance.
@@ 5 Experiments, results & discussion
5.1 Experimental datasets
Table 8: Datasets used in our experiments and their characteristics. The Metric(s) column indicates the distance metrics used in the experiments.
ID Dataset Name Instances Features Classes Field Metric(s)
1 gene Gene expressions 801 20531 Unknown Genomics Euclidean, correlation
2 ceramic Ceramic samples 88 17 Unknown Archaeology Euclidean
3 relig. texts Religious texts 589 8266 Unknown Humanities Cosine, Euclidean, Jaccard
4 conf. papers Conference papers 5804 11463 Unknown Computer science Cosine, Euclidean, Jaccard
5 wdbc Breast Cancer Wisconsin Diagnostic 569 30 2 Healthcare Euclidean
6 wisc Breast Cancer Wisconsin Original 699 9 2 Healthcare Euclidean
7 ecoli Ecoli 336 7 8 Biology Euclidean
8 wine Wine 178 13 3 Chemistry Euclidean
9 iono Ionosphere 351 34 2 Physics Euclidean
10 zoo Zoo 101 16 7 Biology Euclidean
11 iris Iris 150 4 3 Botany Euclidean
12 segment Image Segmentation 2310 19 7 Computer science Euclidean
13 dermatology Dermatology 366 34 6 Healthcare Euclidean
14 balance-scale Balance Scale 625 4 3 Psychology Euclidean
15 heart-statlog Statlog (Heart) 270 13 2 Healthcare Euclidean
16 thy New Thyroid 215 5 3 Healthcare Euclidean
We evaluate the upper bound by applying it to a variety of datasets, both synthetic and real-life. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. (For some parts of the experiments we use the Jaccard distance, which means the data is treated in binary form and similarity is based on the amount of shared attributes.)
An essential part of the evaluation is to establish a tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we adopt an empirical approach: we apply several clustering algorithms to each dataset and take the largest observed $\mathrm{ASW}$ as the lower bound. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
Figure 9: Average silhouette width $\mathrm{ASW}$ vs. number of clusters ($K$). Results on the synthetic dataset 400-64-5-6 for $K$-means, PAM (PAMSIL), and Hierarchical clustering with weighted and single linkage. For each method and each chosen $K$, we compute the resulting $\mathrm{ASW}$. The black dashed line marks the theoretical upper bound $\mathrm{UB}(\Delta)=0.376$; the red dashed line marks the upper bound when a minimum cluster size of $80$ is enforced ($0.282$). The maximum observed $\mathrm{ASW}$ occurs at K=5K=5—a partition into five equal-sized clusters of 80 elements—and is achieved by $K$-means, Hierarchical (weighted linkage), and PAMSIL.

### v2

@@ 4 A data-dependent upper bound of the silhouette
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the remaining data points in the same cluster is not less than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning through the $\Lambda$-quotient of point $i$.
@@
for some $\kappa\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
@@ 20  Append 1q1-q to AA;
The workflow for computing the proposed upper bound is illustrated in Fig.4. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the solution space $\min_{I}|C_{I}|\geq\kappa$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound.
Time complexity
The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$, which takes $\mathcal{O}(n^{2})$ time [medoid].
@@
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). We refer to this as the global upper bound. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space. We will refer to $\texttt{UB}_{\kappa}(\Delta)$ as an adjusted upper bound.
@@
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$. Its tightness is data-dependent.
Use cases
In theory, when sharp, the upper bound $\texttt{UB}(\Delta)$ can be used to confirm that an empirical ASW has reached the global optimum. Even when loose, it increases the interpretability of empirical measurements.
In real-world scenarios where the global upper bound appears loose, we speculate that adjusted versions can provide more practical use.
The individual silhouette width upper bounds generated by Algo. 1 provide a neat complement to standard silhouette visualizations, where the point-wise silhouette widths are typically displayed next to each other in a bar graph.
In the Experiments section, we examine these use cases.
Figure 4: Workflow for computing the proposed $\mathrm{ASW}$ upper bound.
4.2 Extension to the macro-averaged silhouette
The standard $\mathrm{ASW}$ can produce misleading results when the clusters vary greatly in size. In [john], John Pavlopoulos, Georgios Vardakas, and Aristidis Likas show that the so-called macro-averaged silhouette is more robust to cluster-size imbalance. For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
$S_{macro}(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{|C_{I}|}\sum_{i\in C_{I}}s(i).$
In practical scenarios where clusters are imbalanced and small groups also matter, this aggregation strategy is more reasonable. In this subsection, we use Proposition 1 to construct an upper bound for this silhouette variant in a constrained solution space.
Consider an empirical clustering result $\widehat{\mathcal{C}}_{K}=\{\widehat{C}_{I}\}_{I=1}^{K}$. We narrow our scope in this context to include all clusterings whose cluster sizes match those of $\widehat{\mathcal{C}}_{K}$. This narrower solution space is still generally astronomically large. Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark. Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$. Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order. With $N_{0}\coloneqq 1$, it follows from the rearrangement inequality that
$S_{macro}(\mathcal{C}_{K},\Delta)$ $\leq\frac{1}{K}\sum_{I=1}^{K}\sum_{i\in C_{I}}\frac{1}{|C_{I}|}\texttt{ub}_{i}$
$\leq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{N_{I}}\sum_{i=\sum_{J=0}^{I-1}N_{J}}^{\sum_{J=1}^{I}N_{J}}\texttt{ub}_{\sigma(i)}.$
The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space. This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated. Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
@@ 5 Experiments, results & discussion
Table 1: Real-world datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
Dataset Title Instances Features Classes Field Metric
Ceramic Chemical Composition of Ceramic Samples 88 17 2 Physics Manhattan
Customers Wholesale customers 378 7 2 Business Cosine
Dermatology Dermatology 366 34 6 Health and Medicine Cosine
Heart Statlog Statlog (Heart) 231 13 2 Health and Medicine Euclidean
Optdigits Optical Recognition of Handwritten Digits 2296 64 10 Computer Science Euclidean
Rna Gene expression cancer RNA-Seq 808 20531 5 Biology Euclidean
Wdbc Wisconsin Diagnostic Breast Cancer 382 30 2 Health and Medicine Euclidean
Wine Wine 178 13 3 Chemistry Manhattan
Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound.
Dataset $K$ ASW $\texttt{UB}(\Delta)$ Worst-case rel. error
400-64-5-6 5 $0.249$ $0.376$ $0.339$
400-64-2-2 2 $0.673$ $0.673$ $0.000$
400-128-7-3 7 $0.522$ $0.566$ $0.078$
1000-300-5-2 5 $0.668$ $0.678$ $0.015$
10000-32-20-2 20 $0.626$ $0.774$ $0.192$
10000-1024-20-4 20 $0.417$ $0.454$ $0.081$
We evaluate the ASW upper bound by applying it to a variety of datasets, both synthetic and real-world. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
An essential part of the evaluation is to establish a nearly tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we use PAMSIL clustering to obtain an empirical $\mathrm{ASW}$. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies. PAMSIL is a suitable algorithm since its objective is to maximize ASW.
Given an empirical ASW value, say $\widehat{S}$, the performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes a clustering that achieves the global maximum. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical ASW is guaranteed to be within $\epsilon$ of the true maximum.